Two 3 sided spinners are made with numbers 1 ,2,3 on each edge, each with an
equal probability of occuring. If the two spinners are spun ,probability of spinning a
Double 3
Any double
A total of five

Answer by math_tutor2020(3817)


1/3 is the probability of landing on "3" since there's one space we want out of 3 spaces total.
(1/3)*(1/3) = 1/9 is the probability of getting two "3"s in a row assuming each spin is independent of each other.

1/9 also represents the probability of getting double "1"s for similar reasoning. Same goes for double "2"s
1/9 + 1/9 + 1/9 = 3/9 = 1/3 is the probability of any double.

Here are the two ways to add to five
2+3 = 5
3+2 = 5
This is out of 3*3 = 9 ways to spin them.
Therefore, 2/9 is the probability of getting a sum of 5 of the two spins.

---------------------------------


Answers:
Double 3 has probability 1/9
Any double has probability 1/3
A total of five has probability 2/9